OpenStudy (shamil98):

Question in comments.. about logs. let :

OpenStudy (shamil98):

\[\large \log_b A =3 ; \log_b C =2 ; \log_b D = 5 \] What is \[\large \log_b \frac{ A^3 D^4 }{ C^2 }\] I just plug in the value right?

OpenStudy (shamil98):

\[\large \log_b \frac{ 3^3 5^4 }{ 2^2 }\] \[\large \log_b =\frac{ 9 \times 625 }{ 4 } = 1406.25\]

terenzreignz (terenzreignz):

Hardly :P

OpenStudy (shamil98):

Oh, what do I do then?

terenzreignz (terenzreignz):

A few properties of logs.... \[\Large \log_b(MN)=\log_b(M)+\log_b(N)\]\[\Large \log_b\left(\frac{M}{N}\right)=\log_b(M) -\log_b(N) \qquad N\ne 0\]\[\Large \log_bM^\color{green}p =\color{green}p\log_b(M)\]

terenzreignz (terenzreignz):

Use these to your advantage, and expand \[\Large \log_b \frac{A^3D^4}{C^2}\] as best you can, and show me what you get :)

terenzreignz (terenzreignz):

Rule of thumb, so you can do it more quickly (when you get the hang of it, of course) If it's in the numerator, add its log, if it's in the denominator, subtract its log... ^That doesn't really make sense from a strict viewpoint, but hey, it worked for me :>

OpenStudy (shamil98):

So.. \[\large (\log_b A^3 - \log_b D^4) - \log_b C^2\]

OpenStudy (shamil98):

oops that should be a plus sign

OpenStudy (shamil98):

in the parantheses.

terenzreignz (terenzreignz):

Okay then ^_^ \[\large (\log_b A^3 \color{red}+\log_b D^4) - \log_b C^2\]

terenzreignz (terenzreignz):

Now, use that third property, the one involving exponents in logs...

terenzreignz (terenzreignz):

\[\Large \log_bM^\color{green}p =\color{green}p\log_b(M)\]

OpenStudy (shamil98):

\[\large 3 \log_b A + 4 \log_b D - 2 \log_b C\]

terenzreignz (terenzreignz):

Yup... NOW, I believe we have specific known value for each of these expressions... \[\large 3 \color{red}{\log_b A }+ 4 \color{blue}{\log_b D }- 2 \color{green}{\log_b C}\] substitute away ^_^

OpenStudy (shamil98):

Oh! \[\large 3(3) + 4(5) - 2(2) = 29 - 4 = 25\]

OpenStudy (shamil98):

Thank you, that was a big help in understanding this :D

terenzreignz (terenzreignz):

...bingo :)

terenzreignz (terenzreignz):

Your first answer was way off, don't you think? XD

OpenStudy (shamil98):

Haha it was lol